• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Fatiha

    Log entry time 16:02:40 on March18,2005

    Entry number 141845

    keyword=Shift Summary

    Shift crew: F. Benmokhtar, R. Gilman, Y. Jiang
    -----------

    10:40 am -TS2 restarted: the neutron detector single rates screen
    went all black. After the reset, we got the control back.

    11:05 am -Asked for a controlled access, Steve Wood went down to fix
    the HVs.
    11:45 Beam back.


    Number of beam trips:
    ---------------------
    18 beam trips from 8:00-9:00 am
    15 beam trips from 9:00-10:00 am
    28 beam trips from 10:00-11:00 am
    no beam until 11:50 am
    20 beam trips from 12:00-1:00 pm
    22 ................. 1:00-2:00 pm
    18 ................. 2:00-3:00 pm
    20 ................. 3:00-4:00 pm



    List of runs:
    -------------

    2764... 200 K
    2765... 200 K
    2766... 93.6 K Stopped (see log entry # 141825)
    2767... test run
    2768... 125.7 K
    2769... 200 K
    2770... 200 K
    2771... 200 K
    2772... 200 K
    2773... 200 K
    2774... goes to next shift.